# Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal

### v1

@@ 3.3 Surprisal-based Refining
3.4 Supervised Fine-tuning
Following the two-stage pruning by our ASAP framework, we construct the final training dataset using the compressed triplets $\{(Q_{i},\mathbf{C}_{i}^{\prime},A_{i})\}_{i=1}^{N}$. For each instance, we concatenate the pruned CoT ($\mathbf{C}_{i}^{\prime}$) and the final answer ($A_{i}$) to form the complete target response, denoted as $R_{i}$. We then fine-tune a large reasoning model on this dataset. The training goal is to minimize the negative log-likelihood of the target response tokens, conditioned on the input question. Formally, the loss is defined as:
where $r_{i,j}$ is the $j$-th token of the target response $R_{i}$, and $\theta$ represents the parameters of the model being fine-tuned. This supervised fine-tuning process effectively distills the knowledge from our pruning framework into the model. By training on these compact and logically salient examples, the model learns to internalize the efficient reasoning patterns.
4 Experiments
Methods HumanEval LiveCodeBench LeetCodeDataset
HE HE+ v1_v3 v4_v5 All
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 73.78 3051 1.16 68.29 3051 1.16 42.16 7088 3.59 25.37 8336 5.15 19.74 8680 4.95
Original 78.05 2973 1.12 75.61 2973 1.12 52.12 6611 3.15 30.97 8289 4.83 25.00 8485 4.72
Selective Context 59.76 2979 1.13 54.88 2979 1.13 30.23 7025 3.75 16.79 8558 5.35 15.79 8461 4.90
LLMLingua-2 71.34 3075 1.19 68.29 3075 1.19 38.89 6953 3.60 22.76 8474 5.31 17.54 8513 4.81
TokenSkip 76.83 2823 1.07 73.78 2823 1.07 32.35 7095 3.85 20.15 8400 5.37 18.42 8503 4.87
SPIRIT 83.54 2764 1.07 75.61 2764 1.07 50.82 6524 3.09 33.58 7892 4.62 25.00 8186 4.45
ASAP 84.15 2464 0.98 78.66 2464 0.98 54.74 5177 2.09 36.19 6035 2.61 27.63 7541 3.48
Table 1: Experimental results of different methods with DeepSeek-R1-Distill-Qwen-7B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds. The best results are highlighted in bold, and the second-best are underlined.
4.1 Experimental Setup
Models and Datasets
All experiments are conducted on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al. 2025), with DeepSeek-R1-Distill-Qwen-7B as the default backbone across all settings. We use the Python subset of the CodeForces-CoTs (Hugging Face 2025) dataset for training. The dataset consists of high-quality Chain-of-Thought (CoT) samples generated by DeepSeek-R1, making it particularly suitable as training data for competitive programming reasoning tasks.
Benchmarks
We evaluate our method on a suite of widely used code generation benchmarks that span a range of difficulty levels:
  • HumanEval (Chen et al. 2021) consists of 164 hand-crafted programming problems. HumanEval+ (Liu et al. 2023) extends HumanEval with approximately 80 times more unique test cases.
  • LiveCodeBench (Jain et al. 2024) is a comprehensive, contamination-free benchmark that continuously aggregates new problems from competitive programming platforms. We use two subsets: v1_v3 (612 problems) and v4_v5 (268 problems).
  • LeetCodeDataset (Xia et al. 2025b) provides a high-quality evaluation set of 228 problems collected from LeetCode. It covers over 90% of Python problems on the platform, with each problem featuring over 100 test cases.
Baselines
We compare our method against a comprehensive set of baselines:
  • Zero-shot: The original model without any task-specific fine-tuning.
  • Original: The model is fine-tuned on the original, uncompressed CoTs from the training data.
  • Selective Context (Li et al. 2023): Identifies and prunes redundant lexical units in the input using self-information, aiming to retain only the most informative tokens.
  • LLMLingua-2 (Pan et al. 2024): Distills GPT-4’s token importance signals into a lightweight Transformer encoder, which is trained as a token classifier. This approach compresses content using the token classifier.
  • TokenSkip (Xia et al. 2025a): Trains a model to selectively skip less important tokens in a CoT, allowing for controllable compression based on token importance scores.
  • SPIRIT (Cui et al. 2025): Identifies critical reasoning steps by measuring perplexity shifts—steps are retained if their removal significantly increases the perplexity of the remaining text.
Except for the zero-shot baseline, all other methods involve fine-tuning the model using CoTs that have been processed according to their strategies.
Metrics
To comprehensively evaluate each method, we report both accuracy and inference efficiency:
  • Accuracy is measured by Pass@1, which denotes the percentage of problems correctly solved on the first attempt.
  • Efficiency is evaluated using two metrics: the average number of generated tokens and the average generation latency measured in seconds per sample.
Implementation Details
All models are fully fine-tuned (i.e., full-parameter fine-tuning) using Unsloth (Daniel Han and team 2023) for efficiency. For inference, we utilize the vLLM (Kwon et al. 2023) engine to enable fast and efficient decoding. To ensure reproducibility, we use greedy decoding and disable prefix caching. The default generation token budget is adjusted based on task difficulty: 6K for HumanEval and HumanEval+, and 10K for LiveCodeBench and LeetCodeDataset. All experiments are conducted on NVIDIA H20 GPUs. Additional implementation details are provided in Appendix A.
4.2 Main Results
The comprehensive results of our method compared to all baselines are presented in Table 1. The results show that the model fine-tuned on CoTs pruned by ASAP consistently achieves the best trade-off between accuracy and efficiency. It outperforms all baselines in terms of accuracy, while generating the fewest tokens and achieving the lowest generation latency across all benchmarks. Statistical analysis using the Wilcoxon signed-rank test confirms that the improvements of our method over all baselines are statistically significant ($p<0.05$).
A deeper analysis of the results reveals a clear distinction between token-level and step-level pruning strategies. Token-level baselines like Selective Context, LLMLingua-2, and TokenSkip exhibit a significant performance degradation compared to the original CoTs. This is because the token removal disrupts the syntactic structure and semantic coherence of the original reasoning steps. Consequently, the fine-tuning data becomes fragmented and grammatically unnatural, making it difficult for the model to learn the intended logical flow of the CoT. Step-level methods, such as SPIRIT, perform significantly better than token-level pruning methods, due to the preservation of sentence-level integrity. While SPIRIT improves efficiency over the Original with comparable accuracy, our method achieves higher efficiency and accuracy at the same time. This improvement is particularly pronounced on the challenging LiveCodeBench v4_v5 benchmark: ASAP reduces the average number of generated tokens by 23.5% (from 7892 to 6035) and lowers generation latency by 43.5% (from 4.62s to 2.61s), while also achieving a 7.8% improvement in accuracy (Pass@1 increases from 33.58% to 36.19%). These results verify our core hypothesis that pruning reasoning chains based on first-token surprisal is more effective than perplexity or token-level pruning.
4.3 Ablation Study
To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants:
  • ASAP w/o Coarse-grained Pruning: This variant skips Stage 1 and applies only the fine-grained surprisal-based pruning directly to the original CoT.
  • ASAP w/o Fine-grained Pruning: This variant applies only the coarse-grained anchor-guided pruning from Stage 1, omitting the surprisal-based refinement in Stage 2.
  • ASAP w/o Any Pruning: Equivalent to the Original baseline, where the model is fine-tuned on the full, uncompressed CoT without any pruning.
Table 2 reports results on LiveCodeBench v4_v5, which is representative of the consistent trends observed across benchmarks. Additional results are included in Appendix B.
The ablation study demonstrates that both pruning stages are essential and mutually complementary for achieving optimal accuracy and efficiency. First, removing the coarse-grained pruning stage (w/o Coarse-grained Pruning) leads to a drop in both accuracy and efficiency. While the accuracy decrease is modest, the generation latency increases by a substantial 76.2% (from 2.61s to 4.60s), underscoring the importance of stage 1 in eliminating redundancy and narrowing the search space for subsequent processing. Second, removing the fine-grained pruning stage (w/o Fine-grained Pruning) results in a significant degradation across all metrics. The accuracy drops by 12.4% (Pass@1 decreases from 36.19% to 31.72%) relative to the ASAP, and efficiency improvements are largely lost. This highlights that Stage 2—our surprisal-based pruning mechanism—is central to precisely identifying and preserving the most critical logical steps. In summary, the two stages operate in synergy: the coarse-grained anchor-guided pruning preliminarily removes redundant steps, while the fine-grained surprisal-based pruning further extracts the logical core of the reasoning trace. This two-stage process is key to ASAP’s strong performance, significantly outperforming any single-stage variant.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 36.19 6035 2.61
w/o Coarse-grained Pruning 35.07 7735 4.60
w/o Fine-grained Pruning 31.72 8061 4.83
w/o Any Pruning 30.97 8289 4.83
Table 2: Ablation study of different pruning strategies for ASAP on LiveCodeBench v4_v5. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
4.4 Impact of Token Budget
To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate). We compare ASAP against the three strong baselines—SPIRIT, Original, and Zero-shot—across all benchmarks, and observe consistent trends. For clarity, we present results on LiveCodeBench v4_v5 under six budget settings ranging from 2K to 12K tokens. Results for other benchmarks and additional efficiency statistics are provided in Appendix C.
As shown in Figure 3, our findings highlight three key advantages of ASAP: First, ASAP consistently outperforms all baselines across all budget settings on the benchmark. This demonstrates the robustness of our method under different resource constraints. Second, ASAP exhibits smooth performance scaling with respect to the token budget. Unlike baseline methods such as Zero-shot, whose accuracy can fluctuate as the budget increases (e.g., Zero-shot shows a decline from 8K to 10K), ASAP reliably improves with more available tokens. This predictable scaling behavior enables users to confidently configure budget limits to meet specific accuracy or latency targets. Third, ASAP achieves superior performance-efficiency trade-offs. For example, ASAP with just an 8K token budget achieves higher accuracy than SPIRIT and Original at a much larger 12K budget. These results further validate the practical utility of ASAP in real-world scenarios with constrained resources and varying performance requirements.
4.5 Training Efficiency
A central motivation for CoT compression is to reduce the computational burden during training and inference, specifically shortening the lengthy reasoning traces typically required by large reasoning models. Fine-tuning on compressed CoTs not only lowers training-time resource consumption but also helps the model internalize efficient reasoning patterns. To this end, we aim to achieve a high compression ratio across all methods while preserving the core logical steps. More details are provided in Appendix A.
To quantify the training efficiency gains, we report two key metrics in Table 3: the average number of tokens per sample and the average training time measured in seconds per step. The results highlight the training efficiency advantage of ASAP. By generating the most compact yet logically rich CoTs, our approach significantly reduces training overhead. Compared to the uncompressed baseline (Original), our method reduces the number of training tokens by 75.6% and shortens training time by 60.7%. These savings substantially exceed those achieved by all other baselines. ASAP not only achieves a more favorable accuracy-efficiency trade-off at inference time, but also enables a more resource-efficient training process, making it a practical and cost-effective solution for real-world deployment.
Methods Tokens Time
Original 13023 80.11
Selective Context 6722 (-48.4%) 63.41 (-20.9%)
LLMLingua-2 6919 (-46.9%) 65.25 (-18.6%)
TokenSkip 9813 (-24.6%) 77.27 (-3.6%)
SPIRIT 6082 (-53.3%) 57.45 (-28.3%)
ASAP 3178 (-75.6%) 31.48 (-60.7%)
Table 3: Training efficiency comparison. We report the average number of tokens per sample and training time measured in seconds per step. Percentages indicate the reduction relative to the Original baseline.
4.6 Generalization to Different Architectures
To validate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three strong baselines: Zero-shot, Original, and SPIRIT. We observe consistent trends across all benchmarks, so for brevity, we present representative results on two key benchmarks: LiveCodeBench v4_v5 and LeetCodeDataset in Table 4. The results across all benchmarks are reported in the Appendix D.
The results on DeepSeek-R1-Distill-Llama-8B are highly consistent with our findings on the DeepSeek-R1-Distill-Qwen-7B, confirming that the effectiveness of ASAP generalizes beyond a specific model family. As shown in the Table 4, ASAP achieves the highest accuracy on both benchmarks, and the improvements in efficiency are even more pronounced. On LiveCodeBench, for instance, ASAP not only surpasses the accuracy of the Original baseline (32.84% vs. 31.34%) but also generates 49.1% fewer tokens and reduces latency by over 3x (from 8.60s to 2.69s). Compared to the Zero-shot baseline, our method reduces token generation by 50.9% and latency by a remarkable 69.8%. This suggests that ASAP is particularly effective in identifying and distilling the core reasoning patterns, validating its robustness and broad applicability for improving reasoning efficiency across different model families.
Methods LiveCodeBench LeetCodeDataset
Zero-shot 25.00 8508 8.90 27.19 8358 8.65
Original 31.34 8202 8.60 26.32 8413 8.85
SPIRIT 30.22 7913 8.45 26.75 8449 8.73
ASAP 32.84 4175 2.69 27.63 3792 2.42
5 Conclusion
In this paper, we addressed the challenge of balancing performance and efficiency in Chain-of-Thought (CoT) reasoning, where longer reasoning traces improve accuracy but introduce significant redundancy and computational overhead. We proposed ASAP, a novel two-stage pruning framework that first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal. Extensive experiments demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art results on multiple code generation benchmarks while substantially reducing both training and inference costs. Although our current work focuses on competitive programming tasks within the code generation domain, we believe the framework is general and broadly applicable.
Appendix A Implementation Details
Software and Hardware
@@ Appendix A Implementation Details
Fine-tuning Configuration
@@
Inference and Evaluation Protocol
All inference benchmarks were run using the vLLM engine with dtype set to bfloat16 and gpu_memory_utilization set to 0.9. To ensure deterministic and reproducible outputs, we set the sampling temperature to 0.0 and set enable_prefix_caching to False.
Baseline Details
@@
Hyperparameters for Our Method
Our method involves several stages. For the LLM-guided Coarse-grained Pruning stage, we employed DeepSeek-V3 for economic reasons. When generating the Direct CoT, we used a deterministic setting (temperature=0.0, top_p=1.0), while for making the final pruning result, we increased exploration (temperature=2.0, top_p=1.0). For Pattern Matching, the similarity threshold $\tau$ was set to 0.6. Finally, during Surprisal-based Fine-grained Pruning, the maximum token budget was set to 4096 to ensure a deep level of compression.
Appendix B Ablation Study of Different Pruning Strategies
To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants: ASAP w/o Coarse-grained Pruning, ASAP w/o Fine-grained Pruning and ASAP w/o Any Pruning. We present results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, and LeetCodeDatsets benchmarks in Table 5, Table 6, Table 7, and Table 8.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 84.15 2464 0.98
w/o Coarse-grained Pruning 82.32 2839 1.10
w/o Fine-grained Pruning 71.34 2897 1.10
w/o Any Pruning 78.05 2973 1.12
Table 5: Ablation study of different pruning strategies for ASAP on HumanEval. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
@@ Appendix B Ablation Study of Different Pruning Strategies
Appendix C Performance under Different Token Budgets
To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate). We compare ASAP with three strong baselines—SPIRIT, Original, and Zero-shot—on HumanEval, LiveCodeBench, and LeetcodeDataset. For HumanEval, we evaluate the performance under four budget settings, ranging from 1K to 6K tokens. For LiveCodeBench and LeetCodeDataset, we evaluate the performance under six budget settings, ranging from 2K to 12K tokens. Results are shown in Table 9, Table 10 and Table 11.
@@ Appendix C Performance under Different Token Budgets
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
1K 10.37 1007 0.28 15.85 983 0.28 11.59 995 0.28 25.00 946 0.27
2K 45.12 1813 0.53 44.51 1702 0.49 49.39 1690 0.49 56.71 1502 0.44
4K 71.95 2561 0.85 67.68 2511 0.82 76.83 2401 0.80 75.00 2116 0.72
6K 73.78 3051 1.16 78.05 2973 1.12 83.54 2764 1.07 84.15 2464 0.98
(a) Results on HumanEval
Budget Zero-Shot Original SPIRIT ASAP
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
@@
(b) Results on HumanEval+
Table 9: Results of different methods under different budgets on HumanEval. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
. Budget Zero-shot Original SPIRIT ASAP Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat 2K 16.50 1966 0.52 17.16 1920 0.51 18.95 1908 0.51 21.57 1833 0.49 4K 32.68 3499 1.06 30.72 3432 1.05 34.80 3370 1.03 34.97 3244 1.00 6K 39.05 4806 1.70 42.65 4673 1.67 43.14 4605 1.64 46.24 4358 1.54 8K 44.28 5903 2.46 47.71 5723 2.43 51.80 5515 2.27 52.61 4919 1.90 10K 42.16 7088 3.59 52.12 6611 3.15 50.82 6524 3.09 54.74 5177 2.09 12K 43.95 7988 5.10 54.41 7473 4.22 51.63 7362 4.09 55.56 5322 2.27
(a) Results on LiveCodeBench v1_v3
@@
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
@@
(b) Results on LiveCodeBench v4_v5
Table 10: Results of different methods under different budgets on LiveCodeBench. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
@@
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat

### v2

@@ Appendix A Algorithms
Appendix B Prompt Templates and Generated Examples
We prompt the LLM to produce a direct thought: a concise, structured, step-by-step explanation that outlines how to derive the answer from the given question. A direct thought always consists of a small number of clear logical steps that directly contribute to the solution, and terminates with a single, explicit final answer. We use the prompt below to elicit such direct thoughts (the model is instructed to return only a detailed step-by-step solution containing only “Step-by-Step Solution" and “Final Answer").
Prompt for Generating the Direct CoT $\mathcal{P}$ Given a question, please tell me how to get this answer step by step. Question: ``` {question} ``` Answer: ```python {answer} ``` Only return a detailed step-by-step solution (containing only “Step-by-Step Solution" and “Final Answer"). The detailed step-by-step solution is:
The following example illustrates a direct thought for a simple programming task: converting an integer into its binary representation. This example demonstrates the defining characteristics of direct CoTs, namely, concise step-by-step reasoning that directly derives the solution, without redundant exploration or verbose commentary.
Example of a Direct CoT Step-by-Step Solution 1. Understand the Problem: The task is to convert a given positive integer n into its binary representation without leading zeros. The constraints are lenient ($n\leq 10^{6}$), so even a straightforward approach is efficient. 2. Input Handling: Read the integer n from standard input. Since the input consists of a single integer, we can directly use input() and convert it to an integer. 3. Binary Conversion: Python’s built-in bin() function converts an integer to its binary string representation (prefixed with 0b). To remove the prefix, slice the string from the third character onward ([2:]). 4. Output the Result: Print the resulting binary string. This ensures no leading zeros are present. 5. Efficiency: The approach is optimal because conversion and slicing both run in $O(\log n)$ time, and memory usage is minimal. Final Code: ```python n = int(input()) print(bin(n)[2:]) ```
Given the direct thought as a reference, this stage instructs the model to prune redundant reasoning from the original chain of thought while preserving the essential logical structure. Specifically, the model is required to 1) remove unnecessary reasoning steps from the original CoT, 2) retain all key supporting content that aligns with the logic of direct CoT, and 3) strictly preserve the original wording and sentence order without introducing new information. This ensures that the compressed reasoning remains faithful to the original thought process while aligning with the concise, goal-oriented structure of the direct CoT. The following detailed prompt is used to elicit such coarse-grained pruning behavior.
Appendix C Implementation Details
Software and Hardware.
@@ Appendix C Implementation Details
Fine-tuning Configuration.
@@
Inference and Evaluation Protocol.
All inference benchmarks were run using the vLLM engine with dtype set to bfloat16 and gpu_memory_utilization set to 0.9. To ensure deterministic and reproducible outputs, we set the sampling temperature to 0.0 and set enable_prefix_caching to False. The default token budget for generation is adjusted based on the task difficulty. Specifically, it is 2K for GSM8K, 4K for MATH500, 6K for HumanEval+, and 10K for AIME24, AIME25, LiveCodeBench, and LeetCodeDataset. Results with other token budget settings are shown in Appendix F.
Baseline Details.
@@
Hyperparameters for Our Method.
Our method involves several stages. For the LLM-guided Coarse-grained Pruning stage, we employed DeepSeek-V3 for economic reasons. When generating the direct thought $\mathcal{P}$, we used a deterministic setting (temperature=0.0, top_p=1.0), while for making the final pruning result, we increased exploration (temperature=1.0, top_p=1.0). For Pattern Matching, the similarity threshold $\tau$ was set to 0.6. Finally, during Surprisal-based Fine-grained Pruning, the maximum token budget was set to 4096 to ensure a deep level of compression.
Appendix D Effect of Different Components.
To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants: ASAP w/o Coarse-grained Pruning, ASAP w/o Fine-grained Pruning, and ASAP w/o Any Pruning. We present results on the HumanEval+, LiveCodeBench v1_v3, and LeetCodeDatsets benchmarks in Table 6, Table 7, and Table 8.
@@
Appendix E Generalization to Different Architectures
To evaluate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three baselines: Zero-shot, Original, and SPIRIT. The results of the code generation task on the HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, and LeetCodeDataset benchmarks are shown in Table 9. The results of the mathematical reasoning task on the GSM8K, MATH500, AIME24, and AIME25 benchmarks are shown in Table 10.
Methods HE+ LCBv1_v3 LCBv4_v5 LCD
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 64.02 3334 1.86 44.12 7162 6.92 25.00 8508 8.90 27.19 8358 8.65
Original 76.22 2978 1.63 52.61 6614 6.16 31.34 8202 8.60 26.32 8413 8.85
SPIRIT 72.56 3159 1.74 52.61 6280 5.84 30.22 7913 8.45 26.75 8449 8.73
ASAP 76.83 2494 1.30 48.86 3605 2.18 32.84 4175 2.69 27.63 3792 2.42
Table 9: Experimental results of different methods on code generation benchmarks with DeepSeek-R1-Distill-Llama-8B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Methods GSM8K MATH500 AIME24 AIME25
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 79.15 1262 0.36 57.20 2612 1.08 33.33 8445 10.42 26.67 8597 10.54
Original 84.91 1310 0.37 63.00 2534 1.01 36.67 8550 10.04 30.00 8268 10.05
SPIRIT 85.67 1256 0.35 62.60 2533 1.01 36.67 8788 10.04 36.67 8094 9.57
ASAP 87.34 768 0.20 66.00 1734 0.65 36.67 5314 6.97 33.33 5348 7.05
Table 10: Experimental results of different methods on mathematical reasoning benchmarks with DeepSeek-R1-Distill-Llama-8B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix F Performance under Different Token Budgets
To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate). We compare ASAP with three strong baselines—SPIRIT, Original, and Zero-shot—on HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, LeetcodeDataset, GSM8K, MATH500, AIME24, and AIME25. For simpler benchmarks (including HumanEval+, GSM8K, and MATH500), we evaluate the performance under four budget settings, ranging from 1K to 6K tokens. For more complex benchmarks (including LiveCodeBench v1_v3, LiveCodeBench v4_v5, LeetcodeDataset, AIME24, and AIME25), we evaluate the performance under six budget settings, ranging from 2K to 12K tokens. Results are shown in Table 11, Table 12, Table 13, Table 14, Table 15, Table 16, Table 17, and Table 18.
@@ Appendix F Performance under Different Token Budgets
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
@@
Table 11: Results of different methods under different budgets on HumanEval+. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
. Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 16.50 1966 0.52 17.16 1920 0.51 18.95 1908 0.51 21.57 1833 0.49 4K 32.68 3499 1.06 30.72 3432 1.05 34.80 3370 1.03 34.97 3244 1.00 6K 39.05 4806 1.70 42.65 4673 1.67 43.14 4605 1.64 46.24 4358 1.54 8K 44.28 5903 2.46 47.71 5723 2.43 51.80 5515 2.27 52.61 4919 1.90 10K 42.16 7088 3.59 52.12 6611 3.15 50.82 6524 3.09 54.74 5177 2.09 12K 43.95 7988 5.10 54.41 7473 4.22 51.63 7362 4.09 55.56 5322 2.27
Table 12: Results of different methods under different budgets on LiveCodeBench v1_v3. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
@@
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
@@
Table 13: Results of different methods under different budgets on LiveCodeBench v4_v5. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
@@
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$